A sample of 8.5 g NH3 on oxidation produces 4.5 g of NO. Calculate the percent yield.
Reaction: 4 NH3 + 5 O2 → 4 NO + 6 H2O
1
Expert's answer
2018-04-05T07:46:10-0400
4NH 3 + 5O 2 → 4NO + 6H 2 O
Percent yield = ·100%
m(NH 3 ) = 8.5 g
n(NH 3 ) = m/M = 8.5g/17g/mol = 0.5 mol
n(NH 3 ) = n(NO) = 0.5 mol
m(NO) = n·M = 0.5mol·30g/mol = 15 g – theoretical yield
Percent yield = ·100% = 30%